# A blueprint for large-scale quantum-network deployments

### v1

@@ III-F Coexistence and integration approach
In this regard, the ETSI GS QKD 015 architecture focuses on the control plane and abstracts other elements. It explicitly includes the QKD modules and the controller-agent assembly, but it does not include other elements that enable them, such as the shared optical network or the IT systems. Furthermore, that specification does not deal explicitly with the FM. Therefore, this work provides a more detailed architecture than the ETSI GS QKD 015 specification.
@@
The design of each node will be the set of components that performs all the network functionalities, both classical and quantum. In quantum cryptography, this involves both the key management functionalities and those that enable the networking. An example of the former can be a hardware security module (HSM) for the secure handling of cryptographic keys or the management of different subscribers. This type of equipment exists in other fields and is subject to highly demanding standards, so security is not compromised [56]. The latter include the QFP components that enable a controlled and managed forwarding of both quantum signals and quantum-distributed resources. Fig. 4 shows a general scheme of the architecture presented in this work, based on a set of components that allows quantum and optical to coexist.
@@
Figure 4: The image shows the set of components that would enable the coexistence of quantum and classical communication technologies in optical networks. It includes all the elements necessary for the provision of the service. It is relevant that the forwarding module (FM) that transports the key has been represented as an infrastructure component. Despite routing not being a key management functionality [56], it has been usually assumed to be part of the LKMS in the QKD literature. Note that the switch depicted at the bottom is a fiber optic switch. A set of them was installed in the co-located domain to enhance the switching capabilities of the optical network.
@@
From top to bottom, the assembly is arranged to enable both hardware and software-based subscriber traffic, and all of them will require quantum-distributed keys to the network. This distribution is made possible by components that control and manage both, the key and infrastructure components of the network. In a complex case, this will be done in an integral way, with a holistic view of the network. This infrastructure can be a heterogeneous set of quantum systems, together with the FMs, as well as the optical network shared by all the other components. With programmatic interfaces, such as those defined by ETSI GS QKD 004 or ETSI GS QKD 018, applications can request enhanced services from the network thanks to an heterogeneous set of resources.
@@
The result is a node ready to integrate an optical network with quantum technologies focusing on cryptographic applications, mainly based on QKD.
@@ IV MadQCI, a blueprint for quantum networking
This section describes MadQCI to serve as a model for other QCI deployments. The fundamentals described so far were taken into account during its design and deployment, so it is a particularisation of those. The aim is to detail the problems that were addressed and the solutions applied to provide a blueprint for any large-scale quantum network deployment.
In particular, this detail concerns the three years in which the testbed was running during the OpenQKD and CIVIQ projects. Its scope and complexity make it possible to provide a detail that had not been achieved so far in the state of the art, as it was the most complex, largest and longest running testbed for quantum communication technologies in Europe. Fig. 5 depicts the layout of MadQCI during the period detailed.
@@
Additionally, the key and network operation and management was performed as much as possible in an SDN-aware manner. The approach followed was based on the general design outlined in section III-F, but devoted to SDN techniques and technologies. The level of integration with each component depended on its level of SDN-awareness.
@@ IV-A1 Classical-specific equipment
REDIMadrid is the research and education network in the region of Madrid and provides connectivity services to several Madrilenian universities and research centres. Its infrastructure is meant for delivering IP-based connectivity services and it is outfitted with production-grade optical network devices—e.g., Ciena 6500 or Adva FSP 3000—. RM facilities are mainly central offices or communication rooms, which fulfilled the security requirements of trusted zones. The network implements path diversity and redundant channels, so that quantum integration could be trialed without risking the service delivery.
@@
(b) Huawei QKD modules, one receiver and one transmitter. Each module consists of a two height units optical box and a one height unit server for the key distillation. The set of Huawei experimental devices were deployed along the two TID and RM domains, as a sub-network, the co-located domain. All network equipment was shared. R&S Ethernet encryptors were deployed at each node of this co-located domain except Quirón.
@@
The premises available in the Telefónica Innovación Digital domain were production central offices. These facilities enabled the daily operation of communication services by connecting the wired and radio subscriber access networks to the backbone transport networks. As such, they were much more regulated critical infrastructures—e.g., taking pictures of the equipment once installed was not authorised—, but suitable for validating quantum-safe communication solutions. This is also why the MadQCI network was designed in this segment from scratch and used dedicated equipment and non-illuminated fibres.
In both network domains, the quantum-communication technologies, both quantum and networking, were installed in regular 19” rack cabinets and had a power supply of 240 V, which fits the industrial standards. At the Norte and Concepción nodes—i.e., central offices of the TID domain—power inverters were used to adapt the direct current typically used in these facilities. In all the premises, there were regular refrigeration conditions. The optical transmission media were standard fibres—e.g., ITU-T G.652—and production-grade connectors were used, such as LC or SC with different polishes. Fig. 6 shows several examples of the deployed quantum and classical equipment.
For delivering the quantum-safe services, both software and hardware resources were deployed. The hardware devices were network encryption equipment consuming the quantum-distributed to cipher the subscriber traffic while performing their network tasks. High-security Rohde & Schwarz SITLine ETH encryptors were deployed for supporting ciphered L2 connectivity services. Analogously, some ADVA 5TCE-PCTN-10GU+AES10G modules were installed to enable L1 encryption [40]. Regarding the software applications, the domain-specific use cases were demonstrated using the same IT systems that ran the SDN-like control and management system, described in the section IV-C.
@@ IV-A2 Quantum-specific equipment
Three different QKD vendors were present in MadQCI, namely Huawei, ID Quantique and Toshiba. An additional, experimental system, from AIT was also installed remotely in the network as a further demonstration of the integration capabilities of the network. The ID Quantique and Toshiba discrete-variable QKD systems were deployed in both domains. The former ones host in the same chassis several blade-like subsystems, such as the quantum devices itself or optical equipment for the channel adding and dropping. The latter ones were designed to be connected in the middle of a fibre pair that connects a facility, in the nearest point to the ducts. It essentially by-passes every incoming or outgoing optical signal, with some losses, except for its own quantum and signaling channels. The co-located set of Huawei continuous-variable QKD systems rely on external optical equipment, shared with the ID Quantique systems.
It should be noted that QKD technology is currently establishing itself in the market with multiple business models, which conditions their integration. Many manufacturers choose point-to-point solutions, such as HEQA, ID Quantique, QTI, ThinkQuantum or Toshiba in DV-QKD or KEEQuant and LuxQuanta in CV-QKD. Each one shows its specificities; for example, the way the Toshiba systems multiplex channels or the HEQA equipment that integrates a receiver and a transmitter to facilitate the key forwarding. Other business models take advantage of the new QKD techniques to offer other competitive advantages, such as Q*Bird, which uses systems based on measurement-device-independent QKD to offer a configuration based on a central hub with multiple emitters arranged in a star configuration. A potential QKD breakthrough will come when miniaturisation in on-chip photonics enables other form factors, such as small form-factor pluggable (SFP) integrations.
Regarding quantum random number generation systems, the integration of Quside systems was tested remotely at MadQCI. In any case, there are also resources from other manufacturers, such as ID Quantique or Quantum-Industries, with different technical approaches. For quantum networking, there are two main ones; those based on PCI cards for IT systems and rack-mounted solutions. In both cases, the delivered entropy rates are similar.
IV-B Quantum and classical coexistence
@@
For achieving the quantum and optical integration, both quantum and classical channels were allocated in the DWDM grid using the most suitable equipment in each case as stated so far. However, in MadQCI several coexistence schemes were tested, both for quantum-classical and quantum-quantum coexistence. Fig. 7 shows these main coexistence schemes implemented in MadQCI.
Figure 7: Main coexistence schemes used in MadQCI. a) Duplex usage of a fibre pair with QKD and classical signals multiplexed in both directions. Several multiplexing stages are used to manage losses, but a single one could be used. The excess loss is compensated with an amplifier at the classical signal input. b) A bidirectional scheme is implemented with a BiDi-multiplexed fibre devoted to classical channels and the other to quantum signals. An additional approach, not depicted here, is using circulators and an unique fibre strand for a bidirectional coexistence [57]. C) Example of a switched QKD network as implemented by Huawei in the co-located domain. On the left, ROADM-type switching is achieved by combining optical fibre switching with multiplexing (enclosed by a dashed line). On the right, FOADM-style switching, where it would be sufficient to transmit on one or the other channel to reach different destinations. Finally, the classic part shows how the wide receive dynamic range of optical transceivers is exploited to pair a couple of transceivers with different channels, which allows better handling of guardbands.
In the RM domain, the legacy production equipment could be used for aggregation and routing of signals to the terminal equipment. Depending on the case, it was necessary to install ad hoc equipment for the final stages—e.g., a demultiplexer for splitting two quantum coexisting channels—. In addition, some QKD systems operated at 1310 nm instead of the 1550 nm window. Regarding the active components, any classical signal recondition was performed after the signal aggregation, e.g., placing the amplifiers near the transceivers. RM operates an IP-based network and the IT systems that supported the use cases were provided with connectivity by RM network.
In the TID domain, the integration could be more flexible, as the equipment was not shared with other transmissions. In these three nodes a higher density of equipment was deployed and coexistence was more intensively tested. Attack or eavesdropping emulators were additionally installed. Several connectivity technologies were also tested: switch-to-switch Ethernet connectivity using colored transceivers, both in the clear and encrypted (R&S encryptors); OTN-like muxponders by Huawei (OTU4 graming); and the multi-frame optical encryption by ADVA modules. On top of these, network logic was configured to ensure visibility between all QKD, IT and encrypting systems using L2/L3 network devices—e.g., ADVA FSP 150-XG304 or Huawei Quidway S6700 Series—.
@@
This network connectivity is fundamental in QKD since, together with the quantum resources allocated in the optical network, the cryptographic key material have to be securely managed and transported, as described in section IV-C. Thus, there are different components that must co-operate for delivering the quantum-enabled service. However, this co-operation must happen in the trusted scenario required to ensure the information-theoretically security. Secure design principles such as economy of mechanism or access control techniques, as L2/L3 slicing, helped to achieve this.
@@
Additionally, a border link was implemented between both domains to enable testing cross-domain use cases. An ID Quantique QKD system was deployed there and IP connectivity between the IT systems of both domains was provided in the last stage of the iteration. Please, note that the co-located domain, namely the Huawei and R&S devices, was deployed in both networks since the very beginning.
Regarding the co-located domain, Huawei implemented a switched QKD network [5]. The objective was to use transmitters and receivers between any node in the network as needed, and with a lower number of modules installed. In addition, this approach makes it possible to avoid the need for trusted nodes in low-loss scenarios, such as metropolitan or access networks with limited reach. For this purpose, an approach closer to the ROADM style was followed.
Each node was equipped with a fibre optic switch between the fixed optical multiplexers which, in this case, allowed channels 34, 37 and 38 to pass through. This performance of the fixed multiplexers combines with that of the fibre optic switches to achieve a cross-connection comparable to that of a network based on ROADMs but without the amplification typical of the WSSs. In addition, CV-QKD systems could tune their lasers to any DWDM channel, so any connection could be established if a sufficiently low-loss optical path was available and, at the nodes in the path, any of the three channels were not being used by another connection.
As a result, this co-located domain was able to test a total of 34 possible routing combinations between the 10 CV-QKD modules installed on 7 nodes.
Table II aggregates the connectivity achieved on each of the links. Thus, they will be available to the operation and management system described in the next section to provide the end-to-end quantum-enhanced services.
@@
TABLE II: Connectivity achieved on each of the network links, so that they can be exploited by the SDN-like operation and management system as explained in the IV-C section. All domains use the same fibres (and usually the same optical equipment). Distillation and synchronisation channels of the QKD systems are included as classical channels, as well as encrypted channels (e.g., in the Quintín-Quijote link there are 6 classical channels, 2 carrying encrypted information, and 2 quantum channels). The co-located domain (Huawei CV-QKD systems) is characterised using the same links although its all-to-all switching capabilities makes the overall results not comparable. In all cases, the contribution of the co-located domain to the amount of quantum channels is a single quantum channel per link. However, due to the switching capabilities implemented by Huawei, it is possible to find up to four CV-QKD channels in a single fibre [5]. The link losses of this co-located domain have been measured by the QKD systems and they include the intra-node losses.
Link Length [km] Loss [dB] Domain Vendor Keyrate [kb/s] Channels Details
Quintín-Quijote 24.1 7.47 RM ID Quantique 1.95 Class.: 6, Quan.: 2, Encryp.: 2
11.8 Co-loc. Huawei 6.2
Quijote-Quirón 24.5 8.72 RM Class.: 4, Quan.: 1, Encryp.: 1 Only deployed at
12.1 Co-loc. Huawei 3.3 co-located domain
Quijote-Quevedo 7.8 4.00 RM ID Quantique 2.04 Class.: 5, Quan.: 2, Encryp.: 1
4.8 Co-loc. Huawei 11
Quevedo-Quijano 33.1 11.80 RM Toshiba 37.21 Class.: 5, Quan.: 1, Encryp.: 0
Quijano-Quinto 1.9 2.0 RM Toshiba 2857.12 Class.: 5, Quan.: 1, Encryp.: 0 5G core simulator
Distrito-Norte 15.0 8.30 TID ID Quantique 1.47 Class.: 3, Quan.: 2, Encryp.: 1
10.2 Co-loc. Huawei 8.4
Distrito-Concepción 15.3 14.3 TID Toshiba 242.33 Class.: 4, Quan.: 2, Encryp.: 1
20.1 Co-loc. Huawei 0.11 via Norte
Concepción-Norte 8.1 2.9 TID Toshiba 1039.98 Class.: 6, Quan.: 2, Encryp.: 2
11.1 Co-loc. Huawei 7.4
Quevedo-Norte 3.9 3.0 RM, TID ID Quantique 1.38 Class.: 4, Quan.: 2, Encryp.: 1 Border link
4.3 Co-loc. Huawei 12
IV-C Management, administration and operation

### v2

@@ III-F Coexistence and integration approach
In this respect, the architecture implied by ETSI GS QKD 015 focuses on the control plane and abstracts other elements. It explicitly includes the QKD modules and the controller-agent assembly, but it does not include other elements that enable them, such as the shared optical network or the IT systems. Furthermore, that specification does not deal explicitly with the FM. Therefore, this work provides a more detailed architecture clues than the ETSI GS QKD 015 specification.
@@
The design of each node will be the set of components that performs all the network functionalities, both classical and quantum. In quantum cryptography, this involves both the key management functionalities and those that enable the networking. An example of the former can be a hardware security module (HSM) for the secure handling of cryptographic keys or the management of different subscribers. This type of equipment exists in other fields and is subject to highly demanding standards, so security should not be compromised [59]. The latter includes the QFP components that enable a controlled and managed forwarding of both quantum signals and quantum-distributed resources. Fig. 4 shows a general scheme of the architecture implied in this work, based on a set of components that allows quantum and optical technologies to coexist.
@@
Figure 4: The image shows the set of components that would enable the coexistence of quantum and classical communication technologies in optical networks. It includes all the elements necessary for the provision of the service: 1) the cryptographic applications that will consume the key, 2) the network and key control systems and 3) the network elements (NE) both QKD-specific and signal-forwarding NEs. These last elements that guide, aggregate or switch signals do not necessarily have to be located at the trusted nodes, since quantum protocols ensure security over the lightpath, but it is desirable to have control over them. Both NEs are called the (quantum) forwarding plane if the network is SD-QKD, as described in the text. It is also relevant that it includes the forwarding module (FM) that transports the key. Despite routing not being a key management functionality [59], it has been usually assumed to be part of the LKMS in the QKD literature.
@@
From top to bottom, the ensemble is arranged to allow both hardware and software-based subscriber traffic, both of which will require quantum-distributed keys. This distribution is made possible by components that simultaneously control and manage the key and infrastructure network elements (NE). In a complex case, this will be done holistically, in a single network control with a comprehensive view of the network, but several control and management systems may be necessary. Additionally, key management systems [59] will be needed to secure the generated keys, if they are not included in other elements.
Finally, at the bottom, the NEs that are the infrastructure for the rest of the components are represented. Those that are specific to the QKD, such as the FMs that carries the key, and those that forward the signals to enable the connectivity of the former —e.g., the shared optical components or the data network that communicates the FMs— are shown separately. QKD modules have been represented as part of both sets, since they perform both QKD functions —e.g., running QKD protocols such as BB84— and network functions: generating the signals adapted to the media or switching their emission frequency to route the signal in a FOADM network are examples.
If the network is SD-QKD, these infrastructure elements would be the programmable ones. With programmable interfaces, such as those defined by ETSI GS QKD 004 or ETSI GS QKD 018, applications can request enhanced services from the network thanks to a heterogeneous set of resources.
@@
The result is a node ready to be integrated an optical network with quantum technologies focusing on cryptographic applications, mainly based on QKD.
@@ IV MadQCI, a blueprint for quantum networking
This section describes MadQCI as a model for the benefit of other QCI deployments. The fundamentals described so far were taken into account during its design and deployment, so it is a specific realization of those. The aim is to detail the problems that were addressed and the solutions applied to provide a blueprint for any large-scale quantum network deployment of the discussed SDN class.
In particular, thes details concern the three years in which the testbed was running during the OpenQKD and CIVIQ projects. Its scope make it possible to describe features that had not been present so far in the state of the art, as it was the most complex, largest and longest running testbed for quantum communication technologies in Europe. Fig. 5 depicts the layout of MadQCI.
@@
Additionally, the key distribution and network operation and management was performed as much as possible in an SDN-aware manner. The approach followed was based on the general design outlined in section III-F. The level of integration with each component depended on its level of SDN-awareness.
@@ IV-A1 Classical-specific equipment
REDIMadrid is the research and education network in the region of Madrid and provides connectivity services to several Madrilenian universities and research centres. Its infrastructure is meant for delivering IP-based connectivity services and it is outfitted with production-grade optical network devices—e.g., Ciena 6500 or Adva FSP 3000—. RM facilities are mainly central offices or communication rooms, which fulfilled the security requirements of trusted zones (i.e. trusted nodes). The network implements path diversity and redundant channels, so that quantum integration could be tested without risking the service delivery.
@@
(b) Huawei QKD modules, one receiver and one transmitter. Each module consists of a two height units optical box and a one height unit server for key distillation. The set of Huawei experimental devices were deployed along the two TID and RM domains, as a sub-network, the co-located domain. All network equipment was shared. R&S Ethernet encryptors were deployed at each node of this co-located domain except Quirón.
@@
The premises available in the Telefónica Innovación Digital domain were production central offices. These facilities enabled the daily operation of communication services by connecting the wired and radio subscriber access networks to the backbone transport networks. As such, they were much more regulated critical infrastructures—e.g., taking pictures of the equipment once installed was not authorised—, suitable for being considered trusted locations (trusted nodes) and therefore for validating QKD-powered, secure communication solutions. This is also why the MadQCI network was designed in this segment from scratch and used dedicated equipment and non-illuminated fibres.
In both network domains, the quantum-communication technologies, both quantum and networking, were installed in regular 19” rack cabinets and had a power supply of 240 V, which fits the industrial standards. At the Norte and Concepción nodes—i.e., central offices of the TID domain—power inverters were used to adapt the direct current typically used in these facilities. In all the premises, there were regular cooling conditions. The optical transmission media were standard fibres—e.g., ITU-T G.652—and production-grade connectors were used, such as LC or SC with different polishes. Fig. 6 shows several examples of the deployed quantum and classical equipment.
For delivering the quantum-safe services, both software and hardware resources were deployed. The hardware devices were network encryption equipment consuming the quantum-distributed key to cipher the subscriber traffic while performing their network tasks. High-security Rohde & Schwarz SITLine ETH encryptors were deployed for supporting ciphered L2 connectivity services. Analogously, some ADVA 5TCE-PCTN-10GU+AES10G modules were installed to enable L1 encryption [42]. Regarding the software applications, the domain-specific use cases were demonstrated using the same IT systems that ran the SDN-inspired control and management system, described in the section IV-C.
@@ IV-A2 Quantum-specific equipment
Three different QKD vendors were present in MadQCI, namely Huawei, ID Quantique and Toshiba. An additional, experimental system, from AIT was also installed remotely in the network as a further demonstration of the integration capabilities of the network. The ID Quantique and Toshiba discrete-variable QKD systems were deployed in both domains. The former ones host, in the same chassis, several blade-like subsystems, such as the quantum devices themselves or optical equipment for the channel adding and dropping. The latter ones were designed to be connected in the middle of a fibre pair that connects a facility, in the nearest point to the ducts. It essentially by-passes every incoming or outgoing optical signal, with some losses, except for its own quantum and signaling channels. The co-located set of Huawei continuous-variable QKD systems [6] rely on external optical equipment, partially shared with the ID Quantique systems and partially standard one but provided by the vendor.
It should be noted that QKD technology is currently establishing itself in the market with multiple business models, which conditions their integration. Many manufacturers choose point-to-point solutions, such as HEQA, ID Quantique, QTI, ThinkQuantum or Toshiba in DV-QKD or KEEQuant and LuxQuanta in CV-QKD. Each one shows its specificities; for example, the way the Toshiba systems multiplex channels or the HEQA equipment that integrates a receiver and a transmitter to facilitate the key forwarding. Other ones take advantage of the new QKD techniques to offer alrernative competitive advantages, such as, e.g. Q*Bird, which uses systems based on measurement-device-independent QKD to offer a configuration based on a central hub with multiple emitters arranged in a star configuration. A potential QKD breakthrough will come when miniaturisation in on-chip photonics would enable other form factors, such as small form-factor pluggable (SFP) integrations and, particularly, lower equipment costs.
A separate discussion deserves quantum random number generation systems, the integration of Quside systems was tested remotely at MadQCI. In any case, there are also this type of resources from other manufacturers, such as ID Quantique or Quantum-Industries, with different technical approaches. For quantum networking, there are two main ones; those based on PCI cards for IT systems and rack-mounted solutions. In both cases, the delivered entropy rates are similar.
IV-B Quantum and classical coexistence and connectivity
@@
This network connectivity is fundamental in QKD since, together with the quantum resources allocated in the optical network, the cryptographic key material has to be securely managed and transported. Thus, there are different components that must co-operate for delivering a quantum-enabled service. However, this co-operation must happen in the trusted scenario required to allow for information-theoretically security. Secure design principles such as economy of mechanism or access control techniques, as L2/L3 slicing, helped to achieve this.
For achieving the quantum and optical integration, both quantum and classical channels were allocated in the DWDM grid using the most suitable equipment in each case as stated so far. However, in MadQCI several coexistence schemes were tested, both for quantum-classical and quantum-quantum coexistence. Fig. 7 shows the main coexistence schemes implemented in MadQCI.
Figure 7: Main coexistence schemes used in MadQCI. a) Duplex usage of a fibre pair with QKD and classical signals multiplexed in both directions. Several multiplexing stages are used to manage losses, but a single one could be used. The excess loss is compensated with an amplifier at the classical signal input. b) A bidirectional scheme is implemented with a BiDi-multiplexed fibre devoted to classical channels and the other to quantum signals. An additional approach, not depicted here, is using circulators and an unique fibre strand for a bidirectional coexistence [60]. C) Example of a switched QKD network as implemented by Huawei in the co-located domain. On the left, ROADM-type switching is achieved by combining optical fibre switching with multiplexing (enclosed by a dashed line). On the right, FOADM-style switching, where it would be sufficient to transmit on one or the other channel to reach different destinations. Finally, the classical part shows how the wide receive dynamic range of optical transceivers is exploited to pair a couple of transceivers with different channels, which allows better handling of guardbands.
In the RM domain, the legacy production equipment could be used for aggregation and routing of signals to the terminal equipment. Depending on the case, it was necessary to install ad hoc equipment for the final stages—e.g., a demultiplexer for splitting two quantum coexisting channels—. In addition, some QKD systems operated at 1310 nm instead of the 1550 nm window. Regarding the active components, any classical signal reconditioning was performed after the signal aggregation, e.g., placing the amplifiers near the transceivers. RM operates an IP-based network and the IT systems that supported the use cases were provided with connectivity by the RM network.
In the TID domain, the integration could be more flexible, as the equipment was not shared with other transmissions. In the three TID nodes a higher density of equipment was deployed and coexistence was more intensively tested. Additionally, attack or eavesdropping emulators were installed. Several connectivity technologies were also tested: switch-to-switch Ethernet connectivity using colored transceivers, both in the clear and encrypted (R&S encryptors); OTN-like muxponders by Huawei (OTU4 graming); and the multi-frame optical encryption by ADVA modules. On top of these, network logic was configured to ensure visibility between all QKD, IT and encrypting systems using L2/L3 network devices—e.g., ADVA FSP 150-XG304 or Huawei Quidway S6700 Series—.
@@
Additionally, a border link was implemented between both domains to enable testing cross-domain use cases. An ID Quantique QKD system was deployed there and IP connectivity between the IT systems of both domains was provided in the last stage of the iteration. Note that the co-located domain, namely the Huawei and R&S devices, was deployed in both networks from the very beginning.
@@
TABLE II: Classical, quantum and encrypted channels that coexisted in the network per link. Distillation and synchronisation channels of the QKD systems are included as classical channels, as well as encrypted channels —e.g., in the Quintín-Quijote link there are 6 classical channels, 2 carrying encrypted information, and 2 quantum channels—. In all cases, the contribution of the co-located domain to the amount of quantum channels is a single quantum channel per link. However, due to the switching capabilities implemented by Huawei, it is possible to find up to four CV-QKD channels in a single fibre [5].
Link
Classical
channels
Quantum
channels
Encrypted
channels
Quintín-Quijote 6 2 2
Quijote-Quirón 4 1 1
Quijote-Quevedo 5 2 1
Quevedo-Quijano 5 1 0
Quijano-Quinto 5 1 0
Distrito-Norte 3 2 1
Distrito-Concepción 4 2 1
Concepción-Norte 6 2 2
Quevedo-Norte 4 2 1
Regarding the co-located domain, Huawei implemented a switched QKD network [5]. The objective was to use transmitters and receivers between any two nodes in the network as needed, allowing for a lower number of modules to be installed. In addition, this approach makes it possible to avoid the need for trusted nodes in relatively low-loss scenarios, such as metropolitan or access networks with limited reach. For this purpose, a design closer to the ROADM style was followed.
Each node was equipped with a fibre optic switch between the fixed optical multiplexers which, in this case, allowed channels 34, 37 and 38 to pass through. This performance of the fixed multiplexers combined with that of the fibre optic switches allows to achieve a cross-connection comparable to that of a network based on ROADMs but without the amplification typical of the WSSs. In addition, the used CV-QKD systems could tune their lasers to any DWDM channel, so any connection could be established if a sufficiently low-loss optical path was available and, at the nodes in the path, any of the three channels was not being used by another connection.
As a result, this co-located domain was able to test a total of 34 possible direct QKD link combinations between the 10 CV-QKD modules installed at 7 nodes.
As a first metric of this integration, the number of classical, encrypted and quantum channels in each of the network links is shown in Table II. Achieving this coexistence is an outcome in itself and is therefore discussed in more detail in the section V.
Figure 8: Circuit diagram with the possible connectivity between the QKD modules in the network. Between each pair of nodes, all circuits are achieved using a single pair of optical fibres —and mostly the same optical equipment—, including both the quantum and the classical signals. Both 1) the point-to-point connectivity of the ID Quantique and Toshiba systems and 2) the switched connectivity achieved by the Huawei systems in the co-located domain are shown. This last set of circuits is shown in the upper half and the best performing connectivity has been chosen from the possibilities —systems can broadcast on several channels, so different performances are obtained on the same path. See [5] for more detail—. Each circuit shown, together with the generated key rate, can be exploited by the SDN-like operation and management system as explained in the IV-C section.
More relevant to this discussion is the connectivity achieved in the network, given that it was achieved with both point-to-point links and a switched approach. Figure 8 shows that connectivity and the achieved keyrate. It is relevant to note how the performance of the network must be evaluated according to the design objective. The point-to-point approach may perform better in terms of generated keyrate, while the switched approach would have worse numbers. However, if the design intention was to maximise connectivity between QKD modules, to satisfy certain requirements or operate the network in other manners, the switched approach achieves higher connectivity. Other performances can be evaluated on a case-by-case basis, such as end-to-end service latency. In all cases, this connectivity and generated key will be available to the operation and management system described in the next section to provide the enhanced end-to-end quantum services.
IV-C Operation, administration and management